OpenStudy (anonymous):

Find the first five terms of the sequence:

OpenStudy (anonymous):

\[a_{n}=\frac{ a _{n-1} }{ 2 }+4 ; a_{1}=16\] @math&ing001

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_n=\frac{1}{2}a_{n-1} +4 }\) \(\large\color{slate}{ a_1=16 }\) (by the way, you can use ~ as a space.) \(\large\color{slate}{ a_2=\frac{1}2(16) +4=12 }\) \(\large\color{slate}{ a_3=\frac{1}2(12) +4=10 }\)

OpenStudy (solomonzelman):

the pattern doesn't seem to be either geometric or arithmetic, what i would do is just to follow what I did for the 4th and the 5tf term

OpenStudy (solomonzelman):

5th*

OpenStudy (solomonzelman):

seems like it is: subtract 4 subtract 2 subtract 1 and on...

OpenStudy (anonymous):

ok thanks :)

OpenStudy (solomonzelman):

yw